Your company has been approached to bid on a contract to sell 5,600 voice recognition (VR) computer keyboards a year for four years. Due to technological improvements, beyond that time they will be outdated and no sales will be possible. The equipment necessary for the production will cost $4.7 million and will be depreciated on a straight-line basis to a zero salvage value. Production will require an investment in net working capital of $475,000 to be returned at the end of the project, and the equipment can be sold for $465,000 at the end of production. Fixed costs are $650,000 per year, and variable costs are $91 per unit. In addition to the contract, you feel your company can sell 14,600, 16,700, 19,500, and 12,000 additional units to companies in other countries over the next four years, respectively, at a price of $202. This price is fixed. The tax rate is 21 percent, and the required return is 12 percent. Additionally, the president of the company will undertake the project only if it has an NPV of $100,000. What bid price should you set for the contract? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

bid price?
Answer #1

Answer:

First let us calculate NPV from additional units:

Contribution per unit = $202 - 91 = $111

Year 0 Cost of equipment Working capital ($4,700,000) ($475,000) Additional Units sold Contribution ($111 * units sold) Fixed

NPV from sale of additional additions = -$1,253,498.29

Hence required NPV from contract = $1,253,498.29 + $100000 =$1353,498.29

Annual cash flow required from contract = $1,253,498.29 / PV of $1 annuity for 4 years at 12%

= $1,353,498.29 / ((1 - 1 /(1+12%) 4 )/12%)

= $445618.25

Annual contribution required from contract = 445618.25 / (1 - 21%) = $564073.73

Bid price = (564073.73/ 5600) + 91 = $184.29

Bid price = $191.73